### v2

@@ Discussion
Here we present a quantum approach to discovering communities in complex networks. We unify the optimization process for binary, weighted, and directed networks and explicitly consider the resolution parameter to consider multiple topological scales present in the networks. Our experiments show that quantum annealing (QA) can be successfully used to maximize the modularity of a network and pinpoint meaningful groups of highly connected nodes. Its performance is robust across multiple experimental conditions and, at least comparable to alternative methods based on classical heuristics. As opposed to earlier attempts, we restrict ourselves to pure QPUs avoiding hybrid solutions that obscure the algorithm results [24, 25] and provide an account of the possible origins of suboptimal performances. While many community detection algorithms rely on maximizing a quality function [41], not all are well-suited for every type of network. Thus, the same could be said about our proposal, which explicitly relies on the existence of hierarchies. While we show how the Hierarchical annealing process performs well in networks without a clear topological organization (Fig. 2, see also Table 1), we note how and why exploiting inherent hierarchies could be beneficial rather than detrimental in the study of real-life complex systems.
Unveiling hierarchies in complex networks with quantum computing
@@
Biological and social networks are dynamic rather than static systems, even if we study their properties at a fixed point in time. Their evolution is thus constrained to the underlying mechanism that governs the interactions between their constituents. Barabasi and Albert [34] showed that preferential attachment, where highly connected nodes attract more links, leads to scale-free networks with a slower decay of highly connected nodes than in random networks. Preferential attachment can lead to structural hierarchies, but the latter is not guaranteed [35, 42]. Furthermore, life processes themselves are hierarchical [43, 44], inspiring specific methods to discover such organization [45]. In neuroscience, brain network structure is influenced by plasticity mechanisms occurring across temporal and spatial scales. Hebbian plasticity, in particular, has been shown to modulate preferential attachment, resulting in scale-free and clustered network structures [46]. It is therefore not unreasonable to assume some underlying hierarchical structure to design the algorithms. In the context of our work, we show how the Hierarchical annealing process performs exquisitely in these kinds of systems even improving over classical alternatives (Fig. 3). Additionally, network size did not significantly affect performance, suggesting the process is robust and scalable. The slightly suboptimal results we observed likely stem from the hierarchical process itself [47], rather than issues with the annealing steps given that we mitigated its variability by using multiple and independent instances of the quantum hierarchical process.
@@
One of the main advantages of the hierarchical approach we studied, is the direct interpretability of each one of the recursive divisions. We can construct dendrograms of the sequence of divisions that yield the highest modularity indices for any given network. However, the correctness of such dendrograms depends entirely on the optimality of the represented solution. For that, it remains essential to compare the modularity of the community structure with solutions from other algorithms. As expected, the obtained using H. annealing for all the networks studied in this work were extremely competitive, thus offering insights into the hidden structure of these networks (Figs. S.3 and 6; see also Supplementary Figures). There exist some attempts to reveal hierarchies present in complex networks, most notably relying on bottom-up hierarchical clustering methods [45, 47] and the combination of resolution parameters [48], therefore offering a complementary rather alternative description of fractal-like structure. A few other methods have been developed using the Louvain algorithm to unravel hierarchical relationships within complex systems. For example, each iteration of the Louvain algorithm can be interpreted as a step in a hierarchical process [10, 49]. Alternatively, separate instances of the Louvain algorithm may be used to discover community structures within sub-modules [50]. Instead, our approach considers the whole system at each step inside the hierarchy [3, 28] (see also Eq. 9). Because of that, and contrary to the Louvain solution, this guarantees that every step in the hierarchy is divided taking into account the whole system rather than local sub-modules, a desirable feature for strongly nested systems where our method is expected to outperform several alternatives (Fig. 3; see also Supplementary Figures).
@@
Towards quantum-based multi-resolution community detection methods
@@
The specific case where the resolution parameter $\gamma$ is 1, for which we recover known results in the literature [3], simplifies several expressions throughout this work. However, discovering communities via maximizing the modularity function possesses an inherent limitation where communities smaller than a given size are likely to be omitted [27]. Kumpula, et al., proved that this limit persisted for all $\gamma$ independently of the null model used in the definition of $Q$ [51]. Although the resolution parameter was introduced to weigh the topological scale at which communities ought to be detected [4], it was elegantly shown that it could be employed to overcome the resolution limit [52], albeit disregarding the null model (e.g., $\frac{g_{i}g_{j}}{2m}$). Consequently, approaches that use different definitions for modularity function and the parameter $\gamma$ also exist [53, 54, 47, 55]. However, they still fall within spin-glass systems, therefore being possible to tackle using the same formalism we used.
@@
Nonetheless, Newman and Girvan’s quality index [1] along with the explicit inclusion of a resolution parameter [4], remains, up to date, the most common choice for discovering community structure but, unfortunately, the exact value remains, up to date, an arbitrary choice. In short, large values of the resolution will yield smaller communities, while smaller values of $\gamma$ will result in fewer communities of increased size. Hence, we root the parameter $\gamma$ inside our framework for two main reasons: 1) users need not tune the objective function to maximize, and, 2) it easily adapts to multiresolution methods. We maintain $\gamma$ as a free parameter and we derive all our expressions without specifying its value. Our experiments using multiple values of $\gamma$ within a meaningful range show that the Hierarchical annealing process is robust w.r.t. the chosen resolution; even in networks without inherent hierarchies (Fig. S1). This feature allows us to speculate that QA can effectively detect community structure using multiresolution methods, where $\gamma$ is gradually changed or sampled nonuniformly along the optimization process [56]. Further work should address whether this holds or not, the range for which the annealer cannot coordinate the large number of communities found, and the computational time required to obtain the large number of relatively small-size communities expected for $\gamma>3$.
@@
Quantum processing units (QPUs), regardless of type, are known to struggle with encoding large numbers of variables. In one-hot encoding, this number is of the order $N\times M$ where $N$ is the number of nodes in the graph and $M$ is the number of allowed communities. This is suboptimal because there is no guarantee that the network contains $M$ distinct communities, leading to wasted computational resources [24]. Our approach, by contrast, limits the number of variables to the size of the network. Additionally, our method optimizes systems using discrete rather than binary variables, offering another advantage. Current approaches using one-hot encoding require satisfying a constraint for each variable, $\sum_{j=1,M}x_{ij}-1=0,\ \forall i=1,\ldots,N$, which ensures each node is uniquely assigned to a community. These constraints are typically relaxed by introducing Lagrange $\alpha_{i}$ multipliers in the QUBO formulation. However, determining appropriate values for these multipliers often requires manual tuning or guessing, a process that becomes impractical for weighted networks, where even slight variations can alter the optimal value. While constrained optimization methods can be effective [23, 24], they add a layer of complexity. An alternative approach, discrete quadratic modeling with hybrid quantum-classical methods, offers a promising compromise [25] but sacrifices transparency. Although we employed a recursive process, other combinations of multiple binary divisions could be designed to tackle this and other optimization problems [57], potentially leading to quantum-classical algorithms similar to the Louvain method. For example, in [58], a similar hierarchical approach is used to divide the graph of data points into $2^{n}$ equal parts to achieve load balancing for parallel processing.
@@
A limitation of this work is the increased computational time compared to classical heuristics. However, simulated annealing returns very good results at the cost of increased computational times [5]. Our preliminary experiments suggest a roughly linear relationship between the number of communities detected and the elapsed time. Some observed irregularities in timing may be due to repeated instances and calls to the QPUs. Although we mitigated this by caching precomputed graph embeddings, we were unable to fully address the causes of these longer times. While total times never exceeded one minute per network per run, this remains a limitation for larger, highly segregated networks. Notably, the average time consumed by the QPU per binary division, as reported by the D-Wave portal, was approximately 2 milliseconds, comparable to times for the Louvain and Leiden methods. Thus, with further research, quantum annealing and its combination with algorithmic processes could potentially become competitive in this regard while increasing its accuracy.
@@
Lastly, we did not include signed networks in our formalism, but the extension to them represents a crucial step prior to the general acceptance of QA for the discovery of community structure. Signed networks can contain weighted, negative, and directed networks which make their study an active field of research. Although these networks allow the same spin-glass description used throughout this work [59], some definitions are tailored to each scientific discipline [60, 61, 62]. This calls for specific rather than generic descriptions of the function to optimize, something that we wanted to avoid here. However, it is necessary to assess the performance of QA in these types of systems where non-trivial interactions can emerge (e.g., frustration [63]).
Conclusions
In this work, we use the D-Wave Advantage annealer to maximize the modularity function to discover community structure in complex networks. We demonstrate that hierarchically dividing the network and, at each step, optimizing a binary problem effectively eliminates the need to use one-hot encoding and constraints to maximize Newman’s quality function. As opposed to earlier attempts, the number of communities can be left uncontrolled and found by the algorithm. Moreover, extending Newman’s ideas [3, 28], we analytically show how the function to maximize remains formally independent from the resolution parameter and whether the network is undirected, directed, and/or weighted. We thoroughly conduct experiments to assess whether the underlying topology of the network represents a liability to this nested approach finding that the pure quantum process is stable across topologies and even performs better in scale-free networks, which are ubiquitous in biological systems. Furthermore, the same procedure could, in principle, be applied for multiresolution clustering purposes in any given type of network given that the Hierarchical annealing optimization remains largely equivalent to classical counterparts for multiple values of the resolution parameter. Lastly, the hierarchical annealing process we present allows for interpretable results. When adequately compared to classical alternatives, it permits the exploration of hidden hierarchies within complex systems, potentially indicative of several network dysfunctions.
Acknowledgements

### v3

@@ 4 Discussion
Opposite to earlier attempts, we restrict ourselves to pure QPUs, avoiding hybrid solutions that obscure the algorithm results [Negre2020, Wierzbinski2023] and provide an account of the possible origins of suboptimal performances. While many community detection algorithms rely on maximizing a quality function [Fortunato2010], not all are well-suited for every type of network [Ghasemian2020] and alternative definitions of community structure could be better suited [Rosvall2008, Peixoto2014]. Thus, the same could be said about our proposal, which explicitly relies on the existence of hierarchies. While we show how the Hierarchical annealing process performs well in networks without a clear topological organization, we note how and why exploiting inherent hierarchies could be beneficial rather than detrimental in the study of real-life complex systems.
4.1 Unveiling hierarchies in complex networks with quantum computing
@@
Biological and social networks are dynamic rather than static systems, even if we study their properties at a fixed point in time. Their evolution is thus constrained to the underlying mechanism that governs the interactions between their constituents. Barabasi and Albert [Barabasi1999] showed that preferential attachment, where highly connected nodes attract more links, leads to scale-free networks with a slower decay of highly connected nodes than in random networks. Preferential attachment can lead to structural hierarchies, but the latter is not guaranteed [Holme2002, barabasi2003scale]. Furthermore, life processes themselves are hierarchical [oltvai2002life, alon2003biological], inspiring specific methods to discover such organization [SalesPardo-etal_2007]. In neuroscience, brain network structure is influenced by plasticity mechanisms occurring across temporal and spatial scales. Hebbian plasticity, in particular, has been shown to modulate preferential attachment, resulting in scale-free and clustered network structures [Lynn2024]. It is therefore not unreasonable to assume some underlying hierarchical structure to design the algorithms. In the context of our work, we show how the Hierarchical annealing process performs exquisitely in these kinds of systems, even improving over classical alternatives (Fig. 3). The slightly suboptimal results we observed in some cases likely stem from the hierarchical process itself [Arenas2008], rather than issues with the annealing steps. Basically, in networks where the evidence of a scale-free, or at least, power-law degree distribution is weak [Broido2019], our Hierarchical annealing procedure is not guaranteed to produce the best community structure. Oppositely, hierarchical approaches should be the go-to since they are specially designed to exploit and interpret the results.
@@
One of the main advantages of the hierarchical approach we studied is the direct interpretability of each one of the recursive divisions. We can construct dendrograms of the sequence of divisions that yield the highest modularity indices for any given network. However, the correctness of such dendrograms depends entirely on the optimality of the represented solution. For that, it remains essential to compare the modularity of the community structure with solutions from other algorithms. As expected, the results obtained using H. annealing for all the networks studied in this work were extremely competitive, thus offering insights into the hidden structure of these networks (Fig. 7-8). There exist some attempts to reveal hierarchies present in complex networks, most notably relying on bottom-up hierarchical clustering methods [SalesPardo-etal_2007, Arenas2008] and the combination of resolution parameters [lambiotte2013multi], therefore offering a complementary rather than alternative description of fractal-like structure. A few other methods have been developed using the Louvain algorithm to unravel hierarchical relationships within complex systems. For example, each iteration of the Louvain algorithm can be interpreted as a step in a hierarchical process [Blondel2008, meunier2009hierarchical]. Alternatively, separate instances of the Louvain algorithm may be used to discover community structures within sub-modules [bassett2010efficient]. Instead, our approach considers the whole system at each step inside the hierarchy [Newman2006, Leicht2008] (see also Eq. 9). Because of that, and contrary to the Louvain solution, this guarantees that every step in the hierarchy is divided taking into account the whole system rather than local sub-modules, a desirable feature for strongly nested systems where our method is expected to outperform several alternatives (Fig. 3).
@@
4.2 Towards quantum-based multi-resolution community detection methods
@@
The specific case where the resolution parameter $\gamma$ is 1, for which we recover known results in the literature [Newman2006], simplifies several expressions throughout this work. However, discovering communities via maximizing the modularity function possesses an inherent limitation where communities smaller than a given size are likely to be omitted [Fortunato2007]. Kumpula, et al., proved that this limit persisted for all $\gamma$ independently of the null model used in the definition of $Q$ [Kumpula2007]. Although the resolution parameter was introduced to weigh the topological scale at which communities ought to be detected [Reichardt2006], it was elegantly shown that it could be employed to overcome the resolution limit [Traag2011], albeit disregarding the null model (e.g., $\frac{g_{i}g_{j}}{2m}$). Consequently, approaches that use different definitions for the modularity function and the parameter $\gamma$ also exist [reichardt2004detecting, raghavan2007near, Arenas2008, ronhovde2010local]. However, they still fall within spin-glass systems; therefore, it is possible to tackle them using the same formalism we used.
@@
Nonetheless, Newman and Girvan’s quality index [Newman-Girvan_2004], along with the explicit inclusion of a resolution parameter [Reichardt2006], remains, up to date, the most common choice for discovering community structure. Unfortunately, the exact value remains an arbitrary choice. In short, large values of the resolution will yield smaller communities, while smaller values of $\gamma$ will result in fewer communities of increased size. Hence, we rooted the parameter $\gamma$ inside our framework for two main reasons: 1) users need not tune the objective function to maximize, and 2) it easily adapts to multiresolution methods. We maintained $\gamma$ as a free parameter, and we derived all our expressions without specifying its value. This feature allows us to speculate that QA can effectively detect community structure using multiresolution methods, where $\gamma$ is gradually changed or sampled nonuniformly along the optimization process [Jeub2018]. Further work should address whether this holds or not, the range for which the annealer cannot coordinate the large number of communities found, and the computational time required to obtain the large number of relatively small-sized communities expected for $\gamma>3$.
Our experiments using multiple values of $\gamma$ within a meaningful range show that the Hierarchical annealing process is robust even in networks without inherent hierarchies (Fig. S10-S13). Interestingly, in directed scale-free networks, the hierarchical annealing algorithm yielded sub-optimal values for both low and high values of the resolution parameter but remained competitive in the neighborhood of $\gamma=1$ (Fig. S14). Given the stable results for directed networks at $\gamma=1$ (Fig.2d) and for undirected networks across arbitrary $\gamma$, the symmetrization in Eq.13 might distort the topology and hierarchy of the symmetrized network relative to the original directed system. This effect is most pronounced at resolutions that favor either very large or very small communities, suggesting that a dedicated QA-based multi-resolution framework for directed networks may not be directly achievable through symmetric recursive approaches [Leicht2008]. Therefore, although $\gamma$ integrates naturally into the formalism, its value introduces subtle performance dependencies, possibly linked to tree depth, QUBO coefficients, or the physical requirements to embed a given problem, warranting further investigation.
4.3 Alternative encoding strategies
@@
Quantum processing units (QPUs), regardless of type, are known to struggle with encoding large numbers of variables. In one-hot encoding, this number is of the order $N\times M$, where $N$ is the number of nodes in the graph and $M$ is the number of allowed communities. This is suboptimal because there is no guarantee that the network contains $M$ distinct communities, leading to wasted computational resources [Negre2020]. Our approach, by contrast, limits the number of variables to the size of the network. Additionally, our method optimizes systems using discrete rather than binary variables, offering another advantage. Current approaches using one-hot encoding require satisfying a constraint for each variable, $\sum_{j=1,M}x_{ij}-1=0,\ \forall i=1,\ldots,N$, which ensures each node is uniquely assigned to a community. These constraints are typically relaxed by introducing Lagrange multipliers in the QUBO formulation. However, determining appropriate values for these multipliers often requires manual tuning or guessing, a process that becomes impractical for weighted networks, where even slight variations can alter the optimal value. While constrained optimization methods can be effective [Ushijima-Mwesigwa2017, Negre2020], they add a layer of complexity. DQM, a popular hybrid model of DWAVE, can solve discrete cases [Wierzbinski2023], but incurs increased computational requirements (Table S1), sacrifices transparency due to undisclosed classical–quantum allocations, and ultimately exhibits behavior similar to our fully open approach.
While the results presented here were obtained using D-Wave hardware, the hierarchical strategy is broadly applicable to any solver that encodes communities via binary variables. In particular, the method can be implemented on quantum-inspired devices such as the Fujitsu Digital Annealer [jiang_benchmarks_2024] or on emerging quantum annealing systems such as those developed by the Nippon Electric Company (NEC) corporation [yamaji_development_2022]. In the case of gate-based devices, our solution can be used with the Quantum Approximate Optimization Algorithm (QAOA) algorithm [farhi_quantum_2022]. However, practical experiments in this direction are limited by additional issues, especially the high noise level of those devices and the need to set additional variational hyperparameters.
The applicability of our approach to quantum-inspired algorithms depends on the encoding requirements. For instance, the QIEA-net and iQIEA-net algorithms [yuanyuan_quantum_2018] use integer-type encodings to assign communities to nodes. Consequently, our solution is not directly applicable in this context. Being classical, these algorithms are not bound by the binary encoding constraints of quantum devices. In practice, they achieved lower modularities for the Karate network (Table 2). In contrast, the quantum-inspired simulated bifurcation algorithm [li_community_2025], which correctly resolved the Karate club system, adopts a QUBO formulation with one-hot encoding constraints and Lagrange multipliers. While this approach conflicts with our goal of avoiding explicit constraints and one-hot encodings, it could benefit from recursively encoding binary variables, albeit not being affected by the practical aspects of using penalty weights with real annealers.
Finally, although we employed a recursive process, other combinations of multiple binary divisions could be designed to tackle this and other optimization problems [Nembrini2022], potentially leading to quantum-classical algorithms similar to the Louvain method. For example, in [rathore_load_2025], a similar hierarchical approach is used to divide the graph of data points into $2^{n}$ equal parts to achieve load balancing for parallel processing.
4.4 Hardware constraints and scalability
The dynamic range of values in the Ising Hamiltonian and the corresponding QUBO coefficients for optimization problems solved by D-Wave annealing devices may be limited. This arises from the need to represent these values in the D-Wave QPU, which uses an analog process and can result in limited precision for coefficient representation [^2] For details, see the DWAVE documentation on integrated control errors (ICE).. Therefore, if the ratio between the largest and smallest values in QUBO is relatively small, the problem is easier to solve by QPU (Supplementary Material S5; see also Fig. S21). Importantly, the range of the weights considered throughout this work did not compromise performance. Nonetheless, research should focus on mitigating the detrimental impact derived from these misrepresentations. Currently, there is ongoing research to overcome this issue [mucke2025], but no unique or recommended solution exists.
Noise also affects quantum annealers (Supplementary Material S6). An indirect way to assess its presence is through the breaking of chains of physical qubits [Jeong2025], where each chain encodes a single logical qubit. Chains and their breaking, therefore, indicate that the annealer is effectively solving a slightly perturbed version of the original QUBO problem [Grant2022]. Although both noise and chain breaks tended to increase with the number of nodes, they were not associated with a reduced ability to reach good solutions (Figs. S22 and S23). Instead, noise appeared to promote exploration of a broader set of alternatives, some better and some worse than the mean. The modularity $\hat{Q}$ is the best possible solution, not the average, and appropriately setting the chain strength guarantees a competitive performance.
Another key question is whether our QA approach can scale beyond the 166‑node networks tested. Pegasus connectivity limits embeddings to 169 variables on Advantage 5.4 (177 on Advantage 4.1). While heuristic embeddings could occasionally handle slightly larger graphs, our cached clique embeddings proved faster and more reliable, since they are computed once at the start. Consequently, scaling to larger problems necessitates D‑Wave’s hybrid solvers [Wierzbinski2023], but these incur long runtimes and opaque “black‑box” workflows that limit user control. Local compute resources can ease runtime issues, but transparency and flexibility remain unresolved.
In Algorithm 1, while the modularity matrix requires $\mathcal{O}(n^{2})$, the generalized modularity matrix Bkk is computed in linear time, $\mathcal{O}(n)$, since it is a subset of B. Therefore, a realistic estimate of the complexity of the recursion tree is $\mathcal{O}(n\log n)$ corresponding to the average depth of the tree [Newman2006]. The Akra-Bazzi theorem allows for a more formal proof. Highly unbalanced trees have linear complexity, but they are very unlikely. Thus, the overall complexity of the hierarchical procedure is $\mathcal{O}(n^{2}+n\log n+\mathcal{K}_{n})$, where $\mathcal{K}_{n}$ is the complexity attributed to the repeated binary QA steps. Hence, the computation of B dominates over the recursion, regardless of the intermediate quantum-based steps.
The Louvain and Leiden algorithms scale between $\mathcal{O}(n\log n)$ and $\mathcal{O}(n^{2})$ [Lancichinetti2009], placing the classical component of our approach within the range of established methods [Girvan2002, Duch2005, Newman2004]. Annealing time scales inversely with the square of the minimum energy gap, which depends on the Hamiltonian [Farhi2001, Albash2018], yielding $\mathcal{K}_{n}$ complexities from polynomial to exponential. Empirically, QPU access time followed a logarithmic trend, $\mathcal{K}_{n}\sim\mathcal{O}(\log n)$. Although extrapolation to larger networks was not feasible, our results suggest a mixed linear–logarithmic dependence of QPU time on problem size (Table 3). QPU access times remained below 1 s even for the largest tested networks, with a strong dependence on network topology. However, communication with the DWAVE servers, combined with internal processing and queueing, consistently exceeded QPU computation times. This overhead may hinder scaling to larger networks, whereas local QPU access (e.g., in supercomputing centers) or user-specific sessions could substantially improve the availability and responsiveness of quantum resources.
The advantage of Algorithm 1 is most evident for discrete problems, where the one-hot vector dimension $M\leq n$ must accommodate the correct solution. Setting $M\propto n$ increases the annealing complexity to $\mathcal{O}(\mathcal{K}_{n^{2}})$, likely worsened by the need to fit Lagrange multipliers. A more practical estimate for $M$ follows from the resolution limit [Fortunato2007, Kumpula2007], where the maximum number of detectable communities scales as $M\sim\sqrt{n}$, resulting in $\mathcal{O}(\mathcal{K}_{n\sqrt{n}})$. In both cases, nonetheless, recursion is expected to provide a notable speedup while avoiding Lagrange multiplier fitting (Table 4). Consequently, as QPU connectivity and qubit counts improve, hybrid and quantum methods like ours could scale competitively against classical heuristics without compromising the results.
$f(n)$ Recursion complexity One-hot encoding complexity
I. $c$ $O(n\log n)$ $\mathcal{O}(1)$
II. $\log n$ $O(n\log n)$ $\mathcal{O}(\log n)$
III. $n$ $\mathcal{O}(n\log n)$ $\mathcal{O}(n\sqrt{n})$
IV. $n^{s},\ s>1$ $\mathcal{O}(n^{s})$ $\mathcal{O}(n^{s}n^{s/2})$
V. $e^{n^{s}},\ s\in\mathbb{R}^{+}$ $\mathcal{O}(ne^{n^{s}})$ $\mathcal{O}(e^{n^{s}n^{s/2}})$
Table 4: Computational complexities of the recursion in Algorithm 1 and one-hot encoding for different $f(n)$ forms. In the $n\to\infty$ limit, cases III to V favor recursion, whereas cases I and II favor one-hot encoding. The parameter $c$ is any positive and finite constant. The total QA $\mathcal{K}_{n}$ and recursion complexities depend on $f$ via Eq. 18. The corresponding limits and calculations are detailed in the Supplementary Materials (Section S4.3).
@@ V. $e^{n^{s}},\ s\in\mathbb{R}^{+}$ $\mathcal{O}(ne^{n^{s}})$ $\mathcal{O}(e^{n^{s}n^{s/2}})$
We did not study signed networks. Although these networks allow the same spin-glass description used throughout this work [Traag2009], some definitions are tailored to each scientific discipline [fornito2016fundamentals]. This calls for specific rather than generic descriptions of the function to optimize, something that we wanted to avoid here. However, it is necessary to assess the performance of QA in these types of systems where non-trivial interactions can emerge [kirkley2019balance].
5 Conclusions
We use the D-Wave Advantage annealer to maximize modularity via a hierarchical scheme that solves binary subproblems, avoids one-hot encoding, and automatically determines the number of communities. We show analytically that the objective is independent of the resolution parameter and network type. Experiments demonstrate robustness across topologies, with strong performance on scale-free networks, while the hierarchical framework could support multiresolution analysis to yield interpretable community structures.
Acknowledgments